Shift Summary for Day, Jan. 13, 1999


Subject: Shift Summary for Day, Jan. 13, 1999
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:22:57 CST


              Shift Summary for day shift, 13/Jan/99

On Shift: S. Taegar, A. Bellavance

Hours of beam delivered at all intensities: 6.5
Hours of beam delivered at nominal intensity: 6.5
Hours of SHORT_REG data: 6.5
Hours of special runs (laser scans, muon runs...): 0.7 (laser)

Accounting of lost time: TeV heater and vacuum problems; TeV aborts

========================================================================
Runs ended this shift:
Run 15501 - 10.9M events - SHORT_REG832
Run 15502 - LASER_SCAN
Run 15503 - ongoing - SHORT_REG832

Tapes brought to FCC this shift: QKB156-8, QKE074-100

========================================================================
Accesses during shift: none
========================================================================
Overview of shift: datataking for special short regenerator run.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15501 07:00 Run ongoing with occasional HCC ineff. errors
                        all day (ignored).
      83 07:38 --
      149 TeV abort, beam goes away.
              09:12 Ch.13: "Recovery of TeV underway."
              09:23 Run stopped because estimate to beam is at least
                        30 minutes. Flip NM4AN magnet polarity NEG ->
                        POS.
15502 1 09:35 Start long laser scan.
              09:54 Ch.13: "TeV heater replacement underway. Est. 20
                        mins."
              10:17 "Excellent!" Laser scan finishes.
15503 1 10:30 Start short reg. run.
              10:33 Ch.13: "TeV heater replacement underway. Eat. 5
                        mins."
      4 10:36 Beam is back!
      7 Beam gone for one spill.
              10:40 Ch.13: "Linac vacuum problems. Investigating."
      12-14 Beam gone.
      16-19 Beam gone.
              11:00 Ch.13: "Beam to users."
      71-2 Ktevana errors; KFRENDs.
              12:45 Occasional BOA alarms start occuring. MCR is
                        called, but alarms continue. Chambers and
                        livetime still look good. BOA values get up
                        to 1.6.
              14:00 Gas checklist done.
              14:20 MCR called again about BOA. "Experts are looking
                        into it but do not know what is causing it yet."
                        MCR's records show this has been happening on and
                        off since Tuesday.
      186-8 14:41 Switchyard energy conservation in effect. Beam
                        comes back with large periodic structure in
                        chamber currents.
      196 14:54 Chamber current structure is getting better.
      197 Beam abort for one spill.
      205 15:06 Beam abort for one spill.
      207 15:09 Beam abort for one spill.
      214 15:20 Beam abort for one spill.
      217 15:23 Beam abort for one spill.
      221 15:28 Beam abort for one spill.

========================================================================
Items to Follow up on Next Shift

- put MA ADC back into readout at a beam down time or tape change.

- The muon illumination looks funny because the trigger is biased that
way.

- Call Bruce (x4750) for DLT drive repair (ktev1, drive tps2d2) when beam
down for >1 hour, 7am-3pm only.

_______
Angela Bellavance work: (630) 840-2801
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
                                - Arthur C. Clarke



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:22:59 CST